OpenStudy (anonymous):

I'm a little stumped, can anyone help? I will FAN and MEDAL the person who answers this the best. At a fundraiser breakfast, the bill for three glasses of orange juice and five pancake specials is $7.60, whereas the bill for one glass of orange juice and two pancake specials is $2.90 a. Write a system of two equations to model this problem. Let j stand for the price of a glass of juice and let p stand for the price of a pancake special. b. Use the substitution method to solve the system of equations. c. What would be the bill for one glass of orange juice and one pancake special?

OpenStudy (anonymous):

Part A you get: 7.6=3j+5p 2.9=1j+2p

OpenStudy (anonymous):

For B ~ isolate j from the second equation, then substitute it in the first equation

OpenStudy (anonymous):

7.6=3(2.9−2p)+5p

OpenStudy (anonymous):

solve for p can you?

OpenStudy (anonymous):

@Zoe9639 can you solve P?

OpenStudy (anonymous):

I'm still a little confused :/

OpenStudy (anonymous):

Alright, p is 7.6=8.7-1p

OpenStudy (anonymous):

We need to isolate p

OpenStudy (anonymous):

7,6 = 8.7 - p 7.6 - 8.7 = -p -1.1 = -p 1.1 = p

OpenStudy (anonymous):

so the price of one pancake is $1.1

OpenStudy (anonymous):

Thank-you so much!! @AleshGames

OpenStudy (anonymous):

2.9=1j+2p 2.9=1j+2∗1.1 2.9=j+2.2 0.7=j j=0.7

OpenStudy (anonymous):

So the solution to system of equations is : p=1.1, j=0.7

OpenStudy (anonymous):

We done for part B :D

OpenStudy (anonymous):

Need part C?